OpenStudy (anonymous):

I dont know what formula to use: a car traveling initially at 7.0m/s accelerates uniformly at the rate of 0.80m/s^2 for a distance of 245m -what is its velocity at the end of the acceleration? -what is its velocity after it accelerates for 125m -what is its velocity after it accelerates for 67m?

OpenStudy (anonymous):

\[V = u + at\] \[V^2 - u^2 = 2as\]

OpenStudy (anonymous):

V1 = 7m/s A = .8m/s² d = 245 m - v2 = ? v2² = v1² + 2ad

OpenStudy (ash2326):

a car traveling initially at 7.0m/s accelerates uniformly at the rate of 0.80m/s^2 for a distance of 245m v= velocity after traveling a distance s u=initial velocity a=acceleration s=distance traveled \[\large v^2-u^2=2as\] u=7.0 m/s a=0.80 m/s^2 s=245 m Let's plugin the values \[\large v^2-(7)^2=2\times 0.80\times 245\] now find v from this, to find other parts change the value of s